OpenStudy (anonymous):

find the lim as x approaches 25 for sqrt x -5 / x-25 I get the indeterminate form of 0/0 and I know thats wrong But I don't know what to do from there

hartnn (hartnn):

can you rationalize the numerator ?

OpenStudy (anonymous):

the sqrt only covers x

OpenStudy (anonymous):

I think thats what I'm supposed to do but I honestly don't know how

OpenStudy (anonymous):

How do I start?

OpenStudy (kira_yamato):

Use L'hopital's Rule. That's my suggestion.

hartnn (hartnn):

multiply numerator and denominator by the conjugate of sqrt x - 5 which is \(\sqrt x+5 \)

OpenStudy (kira_yamato):

But you'll get a sqrt in the denominator @hartnn

OpenStudy (amistre64):

the bottom is a difference of squares

OpenStudy (anonymous):

ok I am very confused right now.... :/

hartnn (hartnn):

thats ok, and yes, we can factor denominator too as amistre said

OpenStudy (anonymous):

wouldn't a difference of squares work if x^2-25? @amistre64

OpenStudy (kira_yamato):

@hartnn What about L'hopital's?

OpenStudy (amistre64):

any number can be squared ....

OpenStudy (amistre64):

\[(\sqrt{x})^2-5^2\]

hartnn (hartnn):

if L'Hopitals can be used, then its easy, but lets not use it, lets do it without that rule

OpenStudy (anonymous):

so I can put that in the denominator?

OpenStudy (anonymous):

@amistre64

OpenStudy (kira_yamato):

If you say so ^_^

OpenStudy (amistre64):

you can yes, and then you can see that the top and bottom have like factors

OpenStudy (anonymous):

ok @hartnn

OpenStudy (amistre64):

LHop is good, but only if the student has covered it in class. For some reason, teachers fuss at you if you know more than they teach :)

hartnn (hartnn):

i think you guys will handle this, i'll go help others...

OpenStudy (anonymous):

right so there are like factors and I can cancel out

OpenStudy (amistre64):

correct, you cancel out the "hole" and simplify the setup to an equivalent construction

OpenStudy (kira_yamato):

@amistre64 You have a point there... It's weird isn't it... So much of independent learning. That's why I prefer home-schooling...

OpenStudy (amistre64):

Gauss would never have made it in our day and age lol

OpenStudy (kira_yamato):

I agree...

OpenStudy (anonymous):

so would it become sqrt x - 5?

OpenStudy (anonymous):

as in you cancel out the top and you get that in the denominator?

OpenStudy (amistre64):

\[\frac{\sqrt x-5}{x-25}\] \[\frac{\cancel{\sqrt x-5}}{\cancel{(\sqrt x-5)}(\sqrt x+5)}\]

OpenStudy (amistre64):

1/(sqrt(x)+5) if we do it right ;)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

so now would I insert 25?

OpenStudy (anonymous):

or leave it like that?

OpenStudy (amistre64):

id insert the 25 to see what the value of the limit would be.

OpenStudy (amistre64):

"find the lim" assume we want a definite value of some sort, the proces we took just gets us to a point that is not undeterminate is all

OpenStudy (anonymous):

ok so 1/ sqrt 25+5 = 1/5.5 ?

OpenStudy (amistre64):

1/(5+5) is how i see it

OpenStudy (anonymous):

oh your right because the sqrt only covers x

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

so the lim is 1/10

OpenStudy (kira_yamato):

Yep

OpenStudy (amistre64):

correct

OpenStudy (kira_yamato):

@amistre64 L'Hopital's a great cheat code to bypass so many things lol

OpenStudy (anonymous):

Thank you guys so much, could I post another question?

OpenStudy (kira_yamato):

Sure

OpenStudy (anonymous):

It's a graph I don't know how to do

OpenStudy (amistre64):

new questons are best posted as a new post

OpenStudy (anonymous):

i will